A transformer consists of 340. primary windings and 980. secondary windings. If the emf of the primary coil is 30.0 V, what is the emf across the secondary coil?
Answer #1


In a transformer

Ep/Es = Np/Ns


given Np = 340

Ns = 980

Ep = 30 V


Es = ?


30/Es = 340/980

Es = 86.5 V